OpenStudy (anonymous):

Shades R Us charges $20 per day to rent a lounge chair and $15 per day to rent an umbrella. Dan and Lisa paid a total of $245 to rent a lounge chair and an umbrella each during their vacation. Lisa rented the chair and umbrella for 1 day less than Dan. The following equation represents this situation, where x represents the number of days Dan rented the lounge chair and umbrella. 20x + 15x + 20(x − 1) + 15(x − 1) = 245 What is the total amount Dan paid to rent the lounge chair and umbrella during his vacation?

OpenStudy (anonymous):

$80 $105 $140 $175

OpenStudy (anonymous):

it must ne 140 dollars

hero (hero):

Hint: 20(x - 1) = 20x - 20 15(x - 1) = 15x - 15

hero (hero):

2nd Hint: 20x + 15x + 20x - 20 + 15x - 15 = 245

OpenStudy (callisto):

3rd Hint: 20x + 15x + 20x + 15x - 20 - 15 = 245 That is x(20+15+20+15) - (20+15) =245 Can you simplify it and solve for x?

OpenStudy (anonymous):

x would be 40.

OpenStudy (callisto):

How did you get your answer?

hero (hero):

Hint: Dan's amount = 20x + 15x Lisa's amount = 20(x-1) + 15(x-1)

hero (hero):

yeah x = 4, sorry

hero (hero):

my bad @Castillo

OpenStudy (callisto):

x(20+15+20+15) - (20+15) =245 70x - 35 =245 Add 35 to both sides and you'll get 70x -35 +35 = 245 +35 70x = 280 Then divide both sides by 70 and you'll get x =4

hero (hero):

@Callisto , that's what I got as well. I just responded too quickly. I was actually watching 24 season 5 and came here suddenly.

OpenStudy (callisto):

@Hero It's okay @ohaidarien So you should put x=4 into 20x + 15x to calculate the amount Dan's need to pay

OpenStudy (anonymous):

Oh, ok. Thanks, guys. I did get 'x = 4', I just mistyped it. My bad. xD Thank you, though! I got the answer of 140. :DD